OpenStudy (anonymous):

solve for x and y 4x-6y=2 5x+3y=1

OpenStudy (anonymous):

x=2/7 and y=-1/7

OpenStudy (anonymous):

using either elimination or substitution method

OpenStudy (anonymous):

i have to show work too

OpenStudy (anonymous):

1. 4x - 6y = 2 x = (2 + 6y)/4 2. replace x in the 2nd equation by the equality ---> (2 + 6y)/4 5x+3y=1 ----> 5(2+6y)/4 + 3y = 1 = 10 + 30y + 12y = 4 = 42y = -6 ---------> y = -6/42 = -1/7 3. Take any equation and plug in y = -1/7 and you should be able to get the value of x.

OpenStudy (anonymous):

thank you